OpenStudy (anonymous):

If a beam of light from a He-Ne laser has wavelength 543 nm in air, what is its wavelength in water? The index of refraction for water is 1.33. 722 nm , 408 nm , 540 nm , or 360 nm ?

OpenStudy (michele_laino):

the light speed v in water is: \[v = \frac{c}{n}\]

OpenStudy (michele_laino):

n is the refractive index of water

OpenStudy (anonymous):

ok! what can we plug in? :/

OpenStudy (michele_laino):

we can write this: \[{\lambda _2}\nu = \frac{{{\lambda _1}\nu }}{n}\]

OpenStudy (michele_laino):

since the frequency is the same inside those two mediums, then we have: \[\Large {\lambda _2} = \frac{{{\lambda _1}}}{n}\] where \lambda_1=543 nm, and \lambda_2 is the requested wavelength

OpenStudy (anonymous):

okay! so we plug in 543/n? what is n?

OpenStudy (michele_laino):

n=1.33

OpenStudy (anonymous):

so we get 408? 408 nm is our solution?

OpenStudy (michele_laino):

\[\Large {\lambda _2} = \frac{{{\lambda _1}}}{n} = \frac{{543}}{{1.33}} = ...\]

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

yay!! thank you:)

OpenStudy (michele_laino):

:)